OpenStudy (anonymous):

What values for q (0 < q < 2p) satisfy the equation? 2 square2 sin q + 2 = 0

OpenStudy (anonymous):

\[a. \pi/4,3\pi/4 B.3\pi/4, 5\pi/4 C. \pi/4,7\pi/4 \]

OpenStudy (anonymous):

\[2\sqrt{2}\sin(x)+2=0\] \[2\sqrt{2}\sin(x)=-2\] \[\sin(x)=-\frac{\sqrt{2}}{2}\]is a start

OpenStudy (anonymous):

how would u finish that

OpenStudy (anonymous):

i would think of an angle (or number) whose sine is \(-\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

Now pick angles in \((0,2\pi)\).